Problem 1

On January 1, 1991, you are considering the purchase of ABC Company's common share. You have collected the following information on ABC Company:

1. Book value for common share on January 1, 1991 is $12 per share.

2. Predicted net income for 1991 is $2 per share and net income is expected to increase by 10% every year until 1995.

3. ABC is expected to pay $0.2 dividends per year from 1991 to 1995.

4. After 1995, abnormal earnings will grow by 5% per year.

5. ABC's beta is estimated to be 1.2. The risk-free rate of return is estimated to be 6.6% and the market risk premium is 7%. You should use the CAPM model to estimate the equity cost of capital for ABC Company.

Required: Using the abnormal earnings formula, compute the estimated equity value per share of ABC Company on January 1, 1991. Your answers must show the steps of computations for any credits
Answer #1

Valuation of Shares as per Abnormal Earnings Formula method

====> Book Value + Perpetual Value of Future Expected Residual Incomes

Workings:

1. Calculation of Equity cost of Capital

Ke = Rf + Beta( RM - Rf )

= 6.6 + 1.2 ( 7 )

= 15%

2. Calc of Perpetual Value of Future Expected Residual Incomes

Year 1991 1992 1993 1994 1995 1996

Begining book value per share (Bt-1)

12 13.8 15.8 18.02 20.48 23.21

Net Income per Share (EPS)

2 2.2 2.42 2.66 2.93
Less: Dividends per share(D) 0.2 0.2 0.2 0.2 0.2
Change in Retained Earnings (EPS - D) 1.8 2 2.22 2.46 2.73

Ending Book Value per share

(Bt-1 +EPS - D)

13.8 15.8 18.02 20.48 23.21
Net Income per Share 2 2.2 2.42 2.66 2.93 3.08
Less: Per Share Equity Charge (rBt-1) 1.8 2.07 2.37 2.70 3.07 3.48
Residual Income (EPS-Equity Charge) 0.2 0.13 0.05 -0.04 -0.14 -0.40
DF @ 15% 0.8696 0.7561 0.6575 0.5717 0.4972 0.4323
PV 0.1739 0.0983 0.0329 -0.0229 -0.0696 -0.1729

Vo = 12+0.0397 = 12.04